Prepare a statement of cash flows using the direct method.

FILMORE COMPANY Comparative Balance Sheets December 31 2013 2014 25,000 41,000 Assets 33,000 14,000 25,000 Cash Accounts receivable Inventory Property, plant, and equipment 73,000 Less: Accumulated depreciation 27,000) Total 23,000 45.000$126,000 78,000 46.000 (24,000) 54,000 135,000 Liabilities and Stockholders Equity Accounts payable Income taxes payable Bonds payable Common stock Retained earnings Total $23,000 26,000 20,000 25,000 41,000 $135,000 $ 46,000 23,000 10,000 25,000 22,000 $126,000 FILMORE COMPANY Income Statement For the Year Ended December 31, 2014 Sales revenue Cost of goods sold Gross profit Selling expenses Administrative expenses Income from operations Interest expense Income before income taxes Income tax expense Net income $295,000 194,000 101,000 $28,000 9,000 37,000 64,000 3,000 61,000 14,000 47,000 Additional data: 1. Depreciation expense was $6,000 4s IwSEun EA 2. Dividends of $28,000 were declared and paid. 3. During the year, equipment was sold for $10,000 cash. This equipment cost $13,000 originally and had accumulated depreciation of $3,000 at the time of sale. Additional equipment was purchased for $8,000 cash. 4. Instructions are a statement of cash flows using the indivest method. (a) Net cashp $8,000
Answer #1

Statement of cash flows

I. Cash flows from operating activities
Cash receipts from customers 286,000
Cash payment to suppliers - 233,000
Cash payment for income tax -11,000
Selling expenses - 22,000
Administrative expenses - 9,000
Net cash flows from operating activities 11,000
II. Cash flows from investing activities
Sale of equipment 10,000
Purchase of equipment - 8,000
Net cash flows from investing activities 2,000
III. Cash flows from financing activities
Dividend paid - 28,000
Issue of bonds 10,000
Interest paid - 3,000
Net cash used in financing activities - 21,000
Net increase in cash (I + II + III) - 8,000
Cash balance, beginning 33,000
Cash balance, ending 25,000
  1. Cash receipts from customers = Revenue from sales - Increase in accounts receivables

= 295,000 - (23,000 - 14,000)

= 295,000 - 9,000

= $286,000

ii. Cash payment to suppliers = Cost of goods sold + Increase in inventory + Decrease in accounts payable

= 194,000 + (41,000 - 25,000) + (46,000 - 23,000)

= 194,000 + 16,000 + 23,000

= $233,000

iii. Cash payment for income tax = Income tax expense - Increase in income tax payable

= 14,000 - (26,000 - 23,000)

= $11,000

Accumulated depreciation Account

Equipment 3,000 Balance b/d 24,000
Balance c/d 27,000 Income summary 6,000
30,000 30,000

Equipment Account

Balance b/d 78,000 Accumulated depreciation 3,000
Cash (purchase) (bal. fig.) 8,000 Cash (sale) 10,000
Balance c/d 73,000
86,000 86,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks